# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ II Parameter definitions and security assumptions
In each round $k$ $(k=1,2,...,N)$ of protocol, Alice (Bob) selects an intensity setting $a_{k}$ ($b_{k}$) $\in A=\{\mu,\nu,\omega\}$ with probability $p_{a}$ ($p_{b}$), a basis $x_{k}^{A}$ ($x_{k}^{B}$) $\in Basis=\{X,Z\}$ with probability $q_{x}^{A}$ ($q_{x}^{B}$), and a uniform raw key bit $r_{k}^{A}$ ($r_{k}^{B}$) $\in\mathbb{Z}_{2}=\{0,1\}$. Without loss of generality, we assume that the intensity settings satisfy the condition $\mu>\nu>\omega\geq 0$. Alice (Bob) encodes one of the four states defined by $x_{k}^{A}$ ($x_{k}^{B}$) and $r_{k}^{A}$ ($r_{k}^{B}$) in a phase-randomized weak coherent source with intensity setting $a_{k}$ ($b_{k}$), then sending it to Charles through the quantum channel. However, the actual mean photon number of each pulse emitted by Alice (Bob) may not correspond to her (his) intensity setting $a_{k}$ ($b_{k}$) due to intensity correlations, denoted as $\alpha_{k}$ ($\beta_{k}$). Let $\vec{a}_{k}=a_{k},a_{k-1},...,a_{1}$ ($\vec{b}_{k}=b_{k},b_{k-1},...,b_{1}$) represent the sequence of Alice’s (Bob’s) intensity settings up to round $k$, where $a_{j}~{}(b_{j})\in A$ for every $j$. Meanwhile, for simplicity, we assume perfect phase randomization, flawless polarization encoding, and no side channels beyond intensity correlations.
@@
Although measurement assumptions of receiver are unnecessary in MDI QKD, the definitions of yield and error probability are essential. In this protocol, Alice and Bob post-select the events where they use the same basis and the relay detects |$\ket{\psi^{-}}$ or |$\ket{\psi^{+}}$ [26]. Thus, based on the MDI QKD protocol and above security assumptions, one can define both the yield and the error probability when the emission photon number of Alice and Bob in round $k$ ($n_{k}^{A},n_{k}^{B}\in\mathbb{N}$) is $n$ and $m$, respectively. For example, the yield in Z basis and the error probability in X basis can be expressed respectively as
@@ IV Key rate simulation
In the simulation, $Q_{ab}$ and $E_{ab}$ depend on Alice and Bob’s intensity settings $a$ and $b$, the total transmittance of the channel and detector $\eta$, the probability of background counts $Y_{0}$, and misalignment errors $e_{d}$ppendix B provides the calculations of $Q_{ab}$ and $E_{ab}$ using these parameters. Consistent with Refs. [35, 36], since we are considering the asymptotic regime, the key rate does not depend on the probabilities of decoy settings and basis choices. Therefore, we setting $p_{\mu}\approx 1$ and $q_{Z}\approx 1$ to maximizes key rate similar to Ref. [36]. The other simulation parameters are utilized from Ref. [42] as follows. The bidirectional error correction efficiency to $f=1.16$, the probability of background counts $Y_{0}=4\times 10^{-8}$, and misalignment errors $e_{d}=0.0108$. For the total transmittance of the channel and detector, $\eta=\eta_{d}10^{-\alpha L/10}$, where $\eta_{d}$ is the detection efficiency of the detector, $\alpha$ is the attenuation coefficient of the channel, and $L$ is the distance from Alice (Bob) to Charles. In the simulation, we set $\eta_{d}=0.53$ and $\alpha=0.2$ dB/km.
@@ V Application
Considering the correlation features of a MDI QKD system in practice, the theoretical model can be applied to the MDI QKD system and illustrates its actual secure key rate. The MDI QKD system we take as an example contains four intensities, $V$, $D_{1}$, $D_{2}$ and $S$, and sends a random number sequence repeatedly for $1.52\times 10^{6}$ times [43]. The detection results from the single-photon detector (SPD) are obtained in the experiment. Without prejudice to generality, we take Alice’s side as an example to analyze its intensity correlations with $\xi=1$ and the impact on the secure key rate. There are the following three steps to extract the correlation parameters from the measurement result.
@@
As shown in Fig. 2a, the intensity distribution of $SS$ can be fitted to a Gaussian distribution $N(0.204,0.056^{2})$, which attributes to random fluctuations in the system. Additionally, as shown in Fig. 2b, the intensity distribution of $S$ can be fitted to a Gaussian distribution $N(0.203,0.072^{2})$, which is considered as the result of both random fluctuations and intensity correlations. Assuming that the effects of random fluctuations and intensity correlations on the emission intensity are independent, the deviation caused solely by intensity correlations follows Gaussian distribution $N(-0.001,0.045^{2})$. This shows that intensity correlations can lead to a decrease in the average intensity by 0.001 and an increase in fluctuations with a standard deviation of 0.045. Using the $3\sigma$ rule of the Gaussian distribution, we can calculate the range of intensity fluctuations caused by intensity correlations as $[-0.136,0.134]$. Since we consider the intensity of the $SS$ as the intensity without intensity correlations, therefore, intensity setting $a=0.204$ and actual intensity $\alpha\in[a-0.136,a+0.134]$. Therefore, based on Assumption 3 in Sec. II, the relative deviation can be calculated as $\delta=\frac{|a-\alpha|}{a}\approx 0.666$. Similar to the simulation in Sec. IV, we assume $p_{\mu}\approx 1$. Therefore, estimating the fluctuation range of the signal state under intensity correlations is sufficient to calculate the secure key rate. When the average photon number distributions are unknown, no positive key rate is generated, whereas, when the average photon number distributions follow Gaussian distribution, and the TG model is used, the key rate is depicted by the red dashed line in Fig. 1, indicating that the key can only be transmitted over approximately 5 km.

### v3

@@ II Parameter definitions and security assumptions
In each round $k$ $(k=1,2,...,N)$ of protocol, Alice (Bob) selects an intensity setting $a_{k}$ ($b_{k}$) $\in A=\{\mu,\nu,\omega\}$ with probability $p_{a}$ ($p_{b}$), a basis $x_{k}^{A}$ ($x_{k}^{B}$) $\in Basis=\{X,Z\}$ with probability $q_{x}^{A}$ ($q_{x}^{B}$), and a uniform raw key bit $r_{k}^{A}$ ($r_{k}^{B}$) $\in\mathbb{Z}_{2}=\{0,1\}$. Without loss of generality, we assume that the intensity settings satisfy the condition $\mu>\nu>\omega\geq 0$. Alice (Bob) encodes one of the four states defined by $x_{k}^{A}$ ($x_{k}^{B}$) and $r_{k}^{A}$ ($r_{k}^{B}$) in a phase-randomized weak coherent source with intensity setting $a_{k}$ ($b_{k}$), then sending it to Charlie through the quantum channel. However, the actual mean photon number of each pulse emitted by Alice (Bob) may not correspond to her (his) intensity setting $a_{k}$ ($b_{k}$) due to intensity correlations, denoted as $\alpha_{k}$ ($\beta_{k}$). Let $\vec{a}_{k}=a_{k},a_{k-1},...,a_{1}$ ($\vec{b}_{k}=b_{k},b_{k-1},...,b_{1}$) represent the sequence of Alice’s (Bob’s) intensity settings up to round $k$, where $a_{j}~{}(b_{j})\in A$ for every $j$. Meanwhile, for simplicity, we assume perfect phase randomization, flawless polarization encoding, and no side channels beyond intensity correlations.
@@
Although measurement assumptions of receiver are unnecessary in MDI QKD, the definitions of yield and error probability are essential. In this protocol, Alice and Bob post-select the events where they use the same basis and Charlie |$\ket{\psi^{-}}$ or |$\ket{\psi^{+}}$ [26]. Thus, based on the MDI QKD protocol and above security assumptions, one can define both the yield and the error probability when the emission photon number of Alice and Bob in round $k$ ($n_{k}^{A},n_{k}^{B}\in\mathbb{N}$) is $n$ and $m$, respectively. For example, the yield in Z basis and the error probability in X basis can be expressed respectively as
@@ IV Key rate simulation
In the simulation, $Q_{ab}$ and $E_{ab}$ depend on Alice and Bob’s intensity settings $a$ and $b$, the total transmittance of the channel and detector $\eta$, the probability of background counts $Y_{0}$, and misalignment errors $e_{d}$ppendix B provides the calculations of $Q_{ab}$ and $E_{ab}$ using these parameters. Consistent with Refs. [35, 36], since we are considering the asymptotic regime, the key rate does not depend on the probabilities of decoy settings and basis choices. Therefore, we setting $p_{\mu}\approx 1$ and $q_{Z}\approx 1$ to maximizes key rate similar to Ref. [36]. The other simulation parameters are utilized from Ref. [42] as follows. The bidirectional error correction efficiency to $f=1.16$, the probability of background counts $Y_{0}=4\times 10^{-8}$, and misalignment errors $e_{d}=0.0108$. For the total transmittance of the channel and detector, $\eta=\eta_{d}10^{-\alpha L/10}$, where $\eta_{d}$ is the detection efficiency of the detector, $\alpha$ is the attenuation coefficient of the channel, and $L$ is the distance from Alice (Bob) to Charlie. In the simulation, we set $\eta_{d}=0.53$ and $\alpha=0.2$ dB/km.
@@ V Application
Considering the correlation features of a MDI QKD system in practice, the theoretical model can be applied to the MDI QKD system and illustrates its actual secure key rate. The MDI QKD system we take as an example contains four intensities, $V$, $D_{1}$, $D_{2}$ and $S$, and sends a random number sequence repeatedly for $1.52\times 10^{6}$ times [43]. The detection results from the single-photon detector (SPD) are obtained in the experiment. Without prejudice to generality, we take Alice’s side with its intensity correlations $\xi=1$ as an example to demonstrate the impact on the secure key rate, which is simple and practical for a MDI QKD system. There are the following three steps to extract the correlation parameters from the measurement result.
@@
As shown in Fig. 2a, the intensity distribution of $SS$ can be fitted to a Gaussian distribution $N(0.204,0.056^{2})$, which attributes to random fluctuations caused by the SPD and other system fluctuations. Additionally, as shown in Fig. 2b, the intensity distribution of $S$ can be fitted to a Gaussian distribution $N(0.203,0.072^{2})$, which is considered as the result of both random fluctuations and intensity correlations. Assuming that the effects of random fluctuations and intensity correlations on the emission intensity are independent, the deviation caused solely by intensity correlations follows Gaussian distribution $N(-0.001,0.045^{2})$. This shows that intensity correlations can lead to a decrease in the average intensity by 0.001 and an increase in fluctuations with a standard deviation of 0.045. Using the $3\sigma$ rule of the Gaussian distribution, we can calculate the range of intensity fluctuations caused by intensity correlations as $[-0.136,0.134]$. Since we consider the intensity of the $SS$ as the intensity without intensity correlations, therefore, intensity setting $a=0.204$ and actual intensity $\alpha\in[a-0.136,a+0.134]$. Therefore, based on Assumption 3 in Sec. II, the relative deviation can be calculated as $\delta=\frac{|a-\alpha|}{a}\approx 0.666$. Similar to the simulation in Sec. IV, we assume $p_{\mu}\approx 1$. Therefore, estimating the fluctuation range of the signal state under intensity correlations is sufficient to calculate the secure key rate. When the average photon number distributions are unknown, no positive key rate is generated, whereas, when the average photon number distributions follow Gaussian distribution, and the TG model is used, the key rate is depicted by the red dashed line in Fig. 1, indicating that the key can only be transmitted over approximately 5 km.